A reaction is carried out where 8.00 g of sodium m
Answer #1

1)molecular weight of NaCl = 58.44 g/mol

58.44 = 1.0 mole

20.0 g = 20 / 58.44 =0 .34 moles

so 0.34 moles NaCl produced.

2) cording to balaced reaction

2 moles of NaCl produced by consuming 1 mole Cl2

0.34 moles NaCl produced by consuming (0.34 x 1.0) / 2= 0.17 moles Cl2

1 mole Cl2 = 71 g

0.17 moles = 0.17 x 71 = 12.07g Cl2